# Optimal Transport for Brain-Image Alignment: Unveiling Redundancy and Synergy in Neural Information Processing

### v1

@@ Abstract
In this paper, we address these issues through optimal transport (OT) and theoretically demonstrate why OT provides a more effective alignment strategy than MSE. Specifically, we construct a transport plan between brain voxel embeddings and image embeddings, enabling more precise matching. By controlling the amount of transport, we mitigate the influence of redundant information. We apply our alignment model directly to the Brain Captioning task by feeding brain signals into a large language model (LLM) instead of images. Our approach achieves state-of-the-art performance across ten evaluation metrics, surpassing the previous best method by an average of 6.11% in single-subject training and 3.81% in cross-subject training. Additionally, we have uncovered several insightful conclusions that align with existing brain research. We unveil the redundancy and synergy of brain information processing through region masking and data dimensionality reduction visualization experiments. We believe our approach paves the way for a more precise understanding of brain signals in the future. The code is available soon.
@@ 4.2 Optimal Transport
[…]
@@
[…]
@@
[…]
@@ 5.1 Experiment Setup
We ran all experiments and visualization in Natural Scenes Dataset (NSD) [2] which contains images from COCO [32] and corresponding fMRI signals recorded in their experiments. Due to spatial distortion and/or head displacement over the course of a scan session, voxels on the edges of the imaged volume may not obtain a full set of data for that session. In pre-processing, such voxels are detected, deemed “invalid”, and are essentially set to 0 for the whole scan session. For the most part, brain voxels of interest are almost always valid. We use the valid voxels as our fMRI part. The dataset spans 8 subjects who were scanned for 30-40 hours (30-40 separate scanning sessions), where each session consisted of viewing 750 images for 3 seconds each. Images were seen 3 times each across the sessions and were unique to each subject, except for a select 1,000 images which were seen by all the subjects. Four subjects (No. 1, 2, 5, and 7) who had completed all the sessions were selected. We choose the test set of subject 1 for evaluation. The model training follows two scenarios: single-subject training (subject 1) and cross-subject training (subjects 1, 2, 5, and 7). In cross-subject training, two subjects are randomly selected in each iteration to compute the OT Loss for parameter updates.
@@ V P F W B Other BLEU1 BLEU2 BLEU3 BLEU4 METEOR ROUGE CIDEr SPICE CLIP-S RefCLIP-S Valid Voxels
Table 2: Region of Interest.
@@ 5.3 Redundant and Synergistic Interactions
As shown in Fig. 3 and 7, different regions carry different semantic information. Depending on the information contained in the images, the activation situation is also different, for instance, all images contain place information, but not necessarily word information. These semantic attributes influence the activation of different brain regions. To further investigate the contribution of each area to the alignment task, we employed gradient perturbation analysis and voxel-based weighting for evaluation. The results indicate that the visual region has the lowest importance, whereas the Place region exhibits the highest importance(See in Fig. 4). This not only corroborates the previous conclusion that the visual cortex primarily processes low-level information but also confirms the ubiquity of Place-related information in images.
We employed PCA [24] and UMAP [38] to visualize embeddings under both single region and multi regions, aiming to explore their differences. The results indicate that in the visualization of the single region, embeddings appear noticeably entangled and interwoven, which may be attributed to redundant interactions. In contrast, under multi-region synergy, the visualization reveals a clear separation between synergistic embeddings and vision embeddings, potentially reflecting the influence of synergistic interactions (See in Fig. 6).
@@ 5.4 Ablation Studies
Keeping the experimental hyper-parameters and the network architecture unchanged, we replaced the loss function with MSE loss. The results demonstrate that in single-subject training (See in Tab. 1, UMBRAE(MSE)-S1*), all evaluation metrics showed improvement, with an average increase of 6.28% per evaluation metric. In cross-subject training (See in Tab. 1, UMBRAE(MSE)*), all metrics also exhibited enhancements, with an average improvement of 4.27% per evaluation metric. Then, we calculate the embeddings of the same data set based on two models for correlation analysis. As shown in Fig. 1, MSE primarily captures point-wise relationships, resulting in strong correlations only along the diagonal, while failing to account for global relationships. OT extends beyond point-wise correspondences by highlighting correlations beyond the diagonal, demonstrating its ability to capture global structural relationships. Thus, OT not only preserves the point-wise alignment inherent in MSE but also considers the overall distribution, leading to a more comprehensive representation.

### v2

@@ Abstract
In this paper, we address these issues through optimal transport (OT) and theoretically demonstrate why OT provides a more effective alignment strategy than MSE. Specifically, we construct a transport plan between brain voxel embeddings and image embeddings, enabling more precise matching. By controlling the amount of transport, we mitigate the influence of redundant information. We apply our alignment model directly to the Brain Captioning task by feeding brain signals into a large language model (LLM) instead of images. Our approach achieves state-of-the-art performance across ten evaluation metrics, surpassing the previous best method by an average of 6.11% in single-subject training and 3.81% in cross-subject training. Additionally, we have uncovered several insightful conclusions that align with existing brain research. We unveil the redundancy and synergy of brain information processing through region masking and data dimensionality reduction visualization experiments. We believe our approach paves the way for a more precise understanding of brain signals in the future. The code is available at https://github.com/NKUShaw/OT-Alignment4brain-to-image.
@@ 4.2 Optimal Transport
$\min_{\gamma\in\Pi(\mu,\nu)}\int_{X\times Y}C(\mathbf{x},\mathbf{y})d\gamma(\mathbf{x},\mathbf{y})$ (6)
@@
$\int_{Y}\gamma(\mathbf{x},\mathbf{y})d\mathbf{y}=\mu(\mathbf{x}),\quad\int_{X}\gamma(\mathbf{x},\mathbf{y})d\mathbf{x}=\nu(\mathbf{y})$ (7)
@@
$W_{p}(\mu,\nu)=(\inf_{\gamma\in\Pi(\mu,\nu)}\int_{X\times Y}\|\mathbf{x}-\mathbf{y}\|^{p}d\gamma(\mathbf{x},\mathbf{y}))^{\frac{1}{p}}$ (8)
@@ 5.1 Experiment Setup
We ran all experiments and visualization in Natural Scenes Dataset (NSD) [2] which contains images from COCO [32] and corresponding fMRI signals recorded. Due to spatial distortion and/or head displacement over the course of a scan session, voxels on the edges of the imaged volume may not obtain a full set of data for that session. In pre-processing, such voxels are detected, deemed “invalid”, and are essentially set to 0 for the whole scan session. For the most part, brain voxels of interest are almost always valid. We use the valid voxels as our fMRI part. The dataset spans 8 subjects who were scanned for 30-40 hours (30-40 separate scanning sessions), where each session consisted of viewing 750 images for 3 seconds each. Images were seen 3 times each across the sessions and were unique to each subject, except for a select 1,000 images which were seen by all the subjects. Four subjects (No. 1, 2, 5, and 7) who had completed all the sessions were selected. We choose the test set of subject 1 for evaluation. The model training follows two scenarios: single-subject training (subject 1) and cross-subject training (subjects 1, 2, 5, and 7). In cross-subject training, two subjects are randomly selected in each iteration to compute the OT Loss for parameter updates.
@@ V P F W B Other BLEU1 BLEU2 BLEU3 BLEU4 METEOR ROUGE CIDEr SPICE CLIP-S RefCLIP-S Valid Voxels
Table 2: Region of Interest. V(ison), (P)lace, (F)ace, W(ord), and B(ody).
@@ 5.3 Redundant and Synergistic Interactions
As shown in Fig. 3 and 7, different regions carry different semantic information. Based on the information of the images, the activation situation is different, for instance, all images contain place information, but not necessarily word information. These semantic attributes influence the activation of different brain regions. To further investigate the contribution of each regions to the alignment task, we employed gradient perturbation analysis and voxel-based weighting for evaluation. The results indicate that the visual region has the lowest importance, whereas the Place region exhibits the highest importance(See in Fig. 4). This not only corroborates the previous conclusion that the visual cortex primarily processes low-level information but also confirms the ubiquity of Place-related information in images.
We employed PCA [24] and UMAP [38] to visualize embeddings under single and multi-region, aiming to explore their differences. The results indicate that in the visualization of the single region, embeddings appear noticeably interwoven, which may be attributed to redundant interactions. In contrast, under multi-region synergy, the visualization reveals a clear separation between synergistic embeddings and vision embeddings, potentially reflecting the influence of synergistic interactions (See in Fig. 6).
@@
Region Methods S-Score(↑) D-B Index(↓)
5 regions PCA -0.0936 12.7383
V-PFBW PCA 0.0261 5.7202
5 regions UMAP -0.0784 9.1067
V-PFBW UMAP 0.1344 3.6206
Table 3: Quantitative analysis of visualization
@@ 5.4 Ablation Studies
Keeping the experimental hyper-parameters and the network architecture unchanged, MSE leads to lower performance compared to OT (See in Tab. 1), where UMBRAE(MSE)-S1* is with an average decrease of 6.28% per evaluation metric, and UMBRAE(MSE)* is with an average decrease of 4.27% per evaluation metric. Then, we calculate the embeddings of the same data set based on two models for correlation analysis. As shown in Fig. 1, MSE primarily captures point-wise relationships, resulting in strong correlations only along the diagonal, while failing to account for global relationships. OT extends beyond point-wise correspondences by highlighting correlations beyond the diagonal, demonstrating its ability to capture global structural relationships. Thus, OT not only preserves the point-wise alignment inherent in MSE but also considers the overall distribution, leading to a more comprehensive representation.
@@ 6 Conclusion
Acknowledgement
This research has been funded in part by the U.S. National Science Foundation grants CRII 2348177.